# Xing Huang a), Feng-Yu Wang a)

### v2

@@ 3. $(3)$
In the following, we first estimate $\|v_{s}^{\varepsilon,1}\|_{\mathbb{B}}$, $\|v_{s}^{\varepsilon,2}\|_{\mathbb{B}}$, $|\tilde{v}_{s}^{\varepsilon}|$, $\|v_{t}^{\varepsilon}-\hat{v}_{t}\big{\|}_{\mathbb{B}}$ and $|\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{t}|$ in (a)-(e), then we prove (3.35) in (f).
(a) For $v_{t}^{\varepsilon,1}$. Let
$V_{N}:=\frac{(N\land\|V\|_{\mathbb{B}})V}{\|V\|_{\mathbb{B}}},\ \ N\geq 1.$
By (3.11), we have
$\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu}V_{N}(\mu_{\varepsilon})-P_{t}^{\mu}V_{N}(\mu)}{\varepsilon}=I_{t}^{V_{N}}(\mu,\phi),$
$P_{t}^{\mu}V_{N}(\mu_{\varepsilon})-P_{t}^{\mu}V_{N}(\mu)=\int_{0}^{\varepsilon}\frac{\text{\rm{d}}}{\text{\rm{d}}r}P_{t}^{\mu}V_{N}(\mu_{r})\text{\rm{d}}r=\int_{0}^{\varepsilon}I_{t}^{V_{N}}(r,\mu,\phi)\text{\rm{d}}r,\ \ N\geq 1.$
Since $(B_{2})$ implies
(3.36) ${\|V_{N}\|_{\mathbb{B}}\leq\|V\|_{\mathbb{B}}\leq c(1+|\cdot|^{k-1})}$
for some constant $c>0$, by (2.14), (3.24), the same argument to (3.26), and the dominated convergence theorem for $N\rightarrow\infty$, we find a constant $c(\mu,\phi)>0$ depending on $\phi,\mu$ such that
@@
(3.37) $\|v_{t}^{\varepsilon,1}\|_{\mathbb{B}}=\frac{\|P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)\|_{\mathbb{B}}}{\varepsilon}\leq\frac{c(\mu,\phi)\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \ \varepsilon\in(0,1],t\in(0,T],$
@@
$\lim_{\varepsilon\downarrow 0}\Big{\|}\frac{P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)}{\varepsilon}-I_{t}^{V}(\mu,\phi)\Big{\|}_{\mathbb{B}}$
$\leq\lim_{N\rightarrow\infty}\lim_{\varepsilon\downarrow 0}\bigg{\{}\|I_{t}^{V-V_{N}}(\mu,\phi)\|_{\mathbb{B}}+\frac{\|P_{t}^{\mu}(V-V_{N})(\mu_{\varepsilon})-P_{t}^{\mu}(V-V_{N})(\mu)\|_{\mathbb{B}}}{\varepsilon}\bigg{\}}$
$\leq\lim_{N\rightarrow\infty}\frac{c\|\phi\|_{L^{k}(\mu)}}{\sqrt{t}}\bigg{(}\int_{\mathbb{R}^{d}}\mathbb{E}[\|(V-V_{N})(X_{t}^{x,\mu})\|_{\mathbb{B}}^{k^{*}}]\mu(\text{\rm{d}}x)\bigg{)}^{\frac{1}{k^{*}}}=0,\ \ t\in(0,T].$
Hence,
$h_{t,1}(\varepsilon):=\big{\{}v_{t}^{\varepsilon,1}-I_{t}^{V}(\mu,\phi)\big{\}}\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}$
satisfies
@@
(3.38) $\lim_{\varepsilon\rightarrow 0}\|h_{t,1}(\varepsilon)\|_{\mathbb{B}}=0,\ \ \sup_{\varepsilon\in(0,1]}\|h_{t,1}(\varepsilon)\|_{\mathbb{B}}\leq 2c(\mu,\phi)\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},\ \ t\in[0,T].$
(b) For $v_{t}^{\varepsilon,2}$ and $v_{t}^{\varepsilon}$. Recall that $X_{t}^{\mu_{\varepsilon},\mu}$ solves (3.2) with initial value $X_{0}^{\varepsilon}$. Let

### v4

@@ 2. $(2)$
In the following, we first estimate $\|v_{t}^{\varepsilon,i}\|_{\infty}$ and $|v_{t}^{\varepsilon,i}-\hat{v}_{t}^{i}|(i=1,2)$ in steps (b)-(c), then verify (3.31) in step (d).
(b) Estimates on $\|v_{t}^{\varepsilon,i}\|_{\infty},i=1,2$ and $\|v_{t}^{\varepsilon,1}-\hat{v}_{t}^{1}\|_{\infty}$.
By (3.6) for $f=U_{t}(y,\nu,\cdot)$ and (3.3), we obtain
$v_{t}^{\varepsilon,1}(y,\nu)=\frac{P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu)}{\varepsilon}$
$=\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}\Big{(}U_{t}\big{(}y,\nu,X_{t}^{x+r\phi(x),\mu}\big{)}-U_{t}\big{(}y,\nu,\tilde{\theta}^{\lambda,\mu}_{t}(x+r\phi(x))\big{)}\Big{)}$
(3.32) ×$\qquad\qquad\qquad\qquad\times\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X_{s}^{x+r\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+r\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)\text{\rm{d}}r$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}\Big{(}U_{t}\big{(}y,\nu,X_{t}^{x+\varepsilon u\phi(x),\mu}\big{)}-U_{t}\big{(}y,\nu,\tilde{\theta}^{\lambda,\mu}_{t}(x+\varepsilon u\phi(x))\big{)}\Big{)}$
×$\qquad\qquad\qquad\qquad\times\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon u\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon u\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)\text{\rm{d}}u,$
where in the last step, we used the integral transform $r=\varepsilon u$. Similar to (3.21), noting that (B) implies $[U_{t}(y,\nu,\cdot)]_{\alpha}\leq 1$, by (2.10) in Lemma 2.3 for $p=2$ and $z=\tilde{\theta}^{\lambda,\mu}_{t}(x+\varepsilon r\phi(x))$, (3.3), (2.18) and (2.1), we find a constant $c(\mu,\phi)>0$ depending on $\phi,\mu$ such that
(3.33) $\begin{split}&\|v_{t}^{\varepsilon,1}\|_{\infty}=\sup_{y,\nu}\frac{|P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu)|}{\varepsilon}\\ &\leq\sup_{y,\nu}\int_{0}^{1}\int_{\mathbb{R}^{d}}\bigg{|}\mathbb{E}\bigg{[}\Big{(}U_{t}\big{(}y,\nu,X_{t}^{x+\varepsilon r\phi(x),\mu}\big{)}-U_{t}\big{(}y,\nu,\tilde{\theta}^{\lambda,\mu}_{t}(x+\varepsilon r\phi(x)))\Big{)}\\ &\qquad\qquad\qquad\qquad\times\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon r\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon r\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\bigg{|}\mu(\text{\rm{d}}x)\text{\rm{d}}r\\ &\leq\frac{c(\mu,\phi)\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \ \varepsilon\in(0,1],t\in(0,T].\end{split}$
This together with (3.29) and (3.21) implies that for a constant $c(\mu,\phi)>0$
$h_{t,1}^{\varepsilon}(y,\nu):=\big{\{}v_{t}^{\varepsilon,1}(y,\nu)-\hat{v}_{t}^{1}(y,\nu)\big{\}}\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}$
satisfies
(3.34) $\lim_{\varepsilon\rightarrow 0}|h_{t,1}^{\varepsilon}(y,\nu)|=0,\ \ \sup_{\varepsilon\in(0,1]}\sup_{y,\nu}|h_{t,1}^{\varepsilon}(y,\nu)|\leq c(\mu,\phi)\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},\ \ t\in(0,T].$
Next, we estimate $\|v_{t}^{\varepsilon,2}\|_{\infty}.$ Recall that $X_{t}^{x+\varepsilon\phi(x),\mu}$ solves (3.2) with initial value $x+\varepsilon\phi(x)$. For any $x\in\mathbb{R}^{d},s,t\in[0,T]$, let
(3.35) $\begin{split}&R_{t}^{\varepsilon,x}:=\text{\rm{e}}^{\int_{0}^{t}\langle\eta_{s}^{\varepsilon,x},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta_{s}^{\varepsilon,x}|^{2}\text{\rm{d}}s},\\ &\eta_{s}^{\varepsilon,x}:=\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})\\ &\qquad\quad\times\big{\{}b_{s}^{(1)}(X_{s}^{x+\varepsilon\phi(x),\mu},P_{s}^{\ast}\mu_{\varepsilon})-b_{s}^{(1)}(X_{s}^{x+\varepsilon\phi(x),\mu},P_{s}^{\ast}\mu)\big{\}}.\end{split}$
By [21, Lemma 3.2], we have
$b_{t}^{(1)}(y,P_{t}^{\ast}\mu_{\varepsilon})-b_{t}^{(1)}(y,P_{t}^{\ast}\mu)$
(3.36) $=\int_{0}^{1}\frac{\text{\rm{d}}}{\text{\rm{d}}r}b_{t}^{(1)}(y,(1-r)P_{t}^{\ast}\mu+rP_{t}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}D^{E}b_{t}^{(1)}(y,(1-r)P_{t}^{\ast}\mu+rP_{t}^{\ast}\mu_{\varepsilon})(z)(P_{t}^{\ast}\mu_{\varepsilon}-P_{t}^{\ast}\mu)(\text{\rm{d}}z)\text{\rm{d}}r.$
Since (B) implies (A), Lemma 2.6 holds so that we find constants $c_{0},c>0$ such that
(3.37) $\begin{split}|\eta_{s}^{\varepsilon,x}|&\leq c_{0}\mathbb{W}_{\alpha}(P_{s}^{\ast}\mu_{\varepsilon},P_{s}^{\ast}\mu)\leq c\varepsilon\|\phi\|_{L^{k}(\mu)}\frac{\alpha(s^{\frac{1}{2}})}{\sqrt{s}},\ \ s\in[0,T],\ \varepsilon\in[0,1],x\in\mathbb{R}^{d}.\end{split}$
Then by Girsanov’s theorem, for any $x\in\mathbb{R}^{d}$,
$W_{t}^{\varepsilon,x}:=W_{t}-\int_{0}^{t}\eta_{s}^{\varepsilon,x}\text{\rm{d}}s,\ \ s\in[0,T]$
is a Brownian motion under $\mathbb{Q}:=R_{T}^{\varepsilon,x}\mathbb{P}$. Reformulate (3.2) with $x+\varepsilon\phi(x)$ replacing $x$ as
$\text{\rm{d}}X_{t}^{x+\varepsilon\phi(x),\mu}$ =$=\Big{\{}b_{t}^{(0)}(X_{t}^{x+\varepsilon\phi(x),\mu})+b_{t}^{(1)}(X_{t}^{x+\varepsilon\phi(x),\mu},P_{t}^{*}\mu)\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x+\varepsilon\phi(x),\mu})\text{\rm{d}}W_{t}$
=$=\Big{\{}b_{t}^{(0)}(X_{t}^{x+\varepsilon\phi(x),\mu})+b_{t}^{(1)}(X_{t}^{x+\varepsilon\phi(x),\mu},P_{t}^{*}\mu_{\varepsilon})\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x+\varepsilon\phi(x),\mu})\text{\rm{d}}W_{t}^{\varepsilon,x},$
$\qquad\ X_{0}^{x+\varepsilon\phi(x),\mu}=x+\varepsilon\phi(x),\ x\in\mathbb{R}^{d}.$
By the weak uniqueness of (3.2) with $\mu=\mu_{\varepsilon},$ we get
$v_{t}^{\varepsilon,2}(y,\nu)$ $=\frac{P_{t}^{\mu_{\varepsilon}}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})}{\varepsilon}$
(3.38) $=\frac{\int_{\mathbb{R}^{d}}[P_{t}^{\mu_{\varepsilon}}U_{t}(y,\nu,\cdot)(x+\varepsilon\phi(x))-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(x+\varepsilon\phi(x))]\mu(\text{\rm{d}}x)}{\varepsilon}$
$=\frac{1}{\varepsilon}\int_{\mathbb{R}^{d}}\mathbb{E}[U_{t}(y,\nu,X_{t}^{x+\varepsilon\phi(x),\mu})(R_{t}^{\varepsilon,x}-1)]\mu(\text{\rm{d}}x),\ \ t\in[0,T].$
By (3.37), for any $p\geq 1$ there exists a constant $c(p,\mu,\phi)>0$ such that
(3.39) $\mathbb{E}[|R_{t}^{\varepsilon,x}-1|^{p}]\leq c(p,\mu,\phi)\varepsilon^{p}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{d}}s\bigg{)}^{\frac{p}{2}},\ \ t\in[0,T],\varepsilon\in[0,1],x\in\mathbb{R}^{d}.$
Again by (2.10) in Lemma 2.3 for $p=2$ and $z=\tilde{\theta}^{\lambda,\mu}_{t}(x+\varepsilon\phi(x))$, (3.3), (3.39), (3.3), (2.18) and (2.1), we find a constant $c_{1}(\mu,\phi)>0$ such that
(3.40) $\|v_{t}^{\varepsilon,2}\|_{\infty}$ $\leq c_{1}(\mu,\phi)\alpha(t^{\frac{1}{2}})\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in[0,T],\varepsilon\in(0,1].$
This together with (3.33) yields that for some constant $c_{2}(\mu,\phi)>0$,
$\|v_{t}^{\varepsilon}\|_{\infty}^{2}$ $\leq 2\|v_{t}^{\varepsilon,1}\|_{\infty}^{2}+2\|v_{t}^{\varepsilon,2}\|_{\infty}^{2}$
$\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\alpha(t^{\frac{1}{2}})\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{d}}s\bigg{)},\ \ t\in(0,T],\varepsilon\in(0,1].$
By the definition of $\alpha$ and (2.41), we find a constant $c_{3}(\mu,\phi)>0$ such that
(3.41) $\|v_{t}^{\varepsilon}\|_{\infty}^{2}\leq c_{3}(\mu,\phi)\frac{\alpha(t^{\frac{1}{2}})^{2}}{t},\ \ t\in(0,T],\ \ \varepsilon\in(0,1].$
(c) Estimate on $\|v_{t}^{\varepsilon,2}-\hat{v}_{t}^{2}\|_{\infty}$. Similarly to (b), we have
(3.42) $\begin{split}&\frac{R_{t}^{\varepsilon,x}-1}{\varepsilon}=\int_{0}^{t}R_{s}^{\varepsilon,x}\big{\langle}\varepsilon^{-1}\eta_{s}^{\varepsilon,x},\text{\rm{d}}W_{s}\big{\rangle}\\ &=\int_{0}^{t}R_{s}^{\varepsilon,x}\Big{\langle}\frac{\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})[b_{s}^{(1)}(\cdot,P_{s}^{*}\mu_{\varepsilon})-b_{s}^{(1)}(\cdot,P_{s}^{*}\mu)](X_{s}^{x+\varepsilon\phi(x),\mu})}{\varepsilon},\ \text{\rm{d}}W_{s}\Big{\rangle}\\ &=h_{t}(\varepsilon,x)+\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})v_{s}^{\varepsilon}(X_{s}^{x,\mu},P_{s}^{\ast}\mu),\ \text{\rm{d}}W_{s}\Big{\rangle},\ \ x\in\mathbb{R}^{d},\end{split}$
where
$h_{t}(\varepsilon,x):=\int_{0}^{t}\Big{\langle}$ $\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})R_{s}^{\varepsilon,x}\frac{[b_{s}^{(1)}(\cdot,P_{s}^{*}\mu_{\varepsilon})-b_{s}^{(1)}(\cdot,P_{s}^{*}\mu)](X_{s}^{x+\varepsilon\phi(x),\mu})}{\varepsilon}$
$\qquad\qquad\quad-\zeta_{s}(X_{s}^{x,\mu})v_{s}^{\varepsilon}(X_{s}^{x,\mu},P_{s}^{\ast}\mu),\ \text{\rm{d}}W_{s}\Big{\rangle},\ \ x\in\mathbb{R}^{d}$
satisfies
(3.43) $\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|h_{t}(\varepsilon,x)|^{2}\Big{]}=0,\ \ x\in\mathbb{R}^{d}.$
Indeed, by (3.3) and the definition of $v_{s}^{\varepsilon}$, we have
$\frac{[b_{s}^{(1)}(\cdot,P_{s}^{\ast}\mu_{\varepsilon})-b_{s}^{(1)}(\cdot,P_{s}^{\ast}\mu)]((X_{s}^{x+\varepsilon\phi(x),\mu})}{\varepsilon}$
$=\frac{1}{\varepsilon}\int_{0}^{1}\int_{\mathbb{R}^{d}}D^{E}b_{s}^{(1)}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})(z)(P_{s}^{\ast}\mu_{\varepsilon}-P_{s}^{\ast}\mu)(\text{\rm{d}}z)\text{\rm{d}}r$
$=\int_{0}^{1}v_{s}^{\varepsilon}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r.$
This together with the BDG inequality implies
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}|h_{t}(\varepsilon,x)|^{2}\Big{]}$
(3.44) $\leq 2\int_{0}^{T}\mathbb{E}\Big{|}\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})R_{s}^{\varepsilon,x}\int_{0}^{1}v_{s}^{\varepsilon}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r$
$\qquad\qquad\quad-\zeta_{s}(X_{s}^{x,\mu})v_{s}^{\varepsilon}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)\Big{|}^{2}\text{\rm{d}}s.$
By (3.3), for any $p>1$, we can find a constant $c_{p}>0$ such that
(3.45) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{x+\varepsilon\phi(x),\mu}-X_{t}^{x,\mu}|^{p}\Big{]}\leq c_{p}|\phi(x)|^{p}\varepsilon^{p},\ \ \varepsilon\in[0,1],\mu\in\mathscr{P}_{k}.$
By the boundedness and continuity of $\zeta$ due to (B), $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t<\infty$, (3.45), (3.39), (3.41), (3.3), and the dominated convergence theorem, to prove (3.43), it is sufficient to prove that for $(s,x,r)\in(0,T]\times\mathbb{R}^{d}\times[0,1]$,
(3.46) $\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{|}v_{s}^{\varepsilon}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})-v_{s}^{\varepsilon}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)\Big{|}=0.$
For any $(\omega,\omega^{\prime})\in\Omega\times\Omega$, let
$U^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})=U_{s}(X_{s}^{x+\varepsilon\phi(x),\mu}(\omega^{\prime}),(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon},X_{s}^{y+\varepsilon u\phi(y),\mu}(\omega)),$
$U^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})=U_{s}(X_{s}^{x+\varepsilon\phi(x),\mu}(\omega^{\prime}),(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon},\tilde{\theta}^{\lambda,\mu}_{s}(y+\varepsilon u\phi(y))),$
$\tilde{U}^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})=U_{s}(X_{s}^{x,\mu}(\omega^{\prime}),P_{s}^{\ast}\mu,X_{s}^{y+\varepsilon u\phi(y),\mu}(\omega)),$
$\tilde{U}^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})=U_{s}(X_{s}^{x,\mu}(\omega^{\prime}),P_{s}^{\ast}\mu,\tilde{\theta}^{\lambda,\mu}_{s}(y+\varepsilon u\phi(y))).$
Since (B) implies (A), (2.29) holds such that
(3.47) $\mathbb{W}_{k}((1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon},P_{s}^{\ast}\mu)\leq r\mathbb{W}_{k}(P_{s}^{\ast}\mu_{\varepsilon},P_{s}^{\ast}\mu)\leq cr\varepsilon\|\phi\|_{L^{k}(\mu)}.$
By (3.3), (3.3) and Hölder’s inequality, we conclude that for any $\beta\in(1,k)$,
$\mathbb{E}|v_{s}^{\varepsilon,1}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})-v_{s}^{\varepsilon,1}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)|$
$\leq\int_{0}^{1}\int_{\mathbb{R}^{d}}\int_{\Omega\times\Omega}\bigg{|}\big{[}(U^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-U^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))$
$\qquad\qquad\qquad\quad-(\tilde{U}^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-\tilde{U}^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))\big{]}$
×$\qquad\qquad\quad\times\int_{0}^{s}\frac{1}{s}\big{\langle}\zeta_{v}(X_{v}^{y+\varepsilon u\phi(y),\mu})\nabla_{\phi(y)}X_{v}^{y+\varepsilon u\phi(y),\mu},\text{\rm{d}}W_{v}\big{\rangle}\bigg{|}\text{\rm{d}}\mathbb{P}(\omega)\text{\rm{d}}\mathbb{P}(\omega^{\prime})\mu(\text{\rm{d}}y)\text{\rm{d}}u$
@@
$\leq c_{0}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{\sqrt{s}}|\phi(y)|\Big{\{}\int_{\Omega\times\Omega}\bigg{|}(U^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-U^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))$
$\qquad\qquad\qquad\quad-(\tilde{U}^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-\tilde{U}^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))\bigg{|}^{\beta}\text{\rm{d}}\mathbb{P}(\omega)\text{\rm{d}}\mathbb{P}(\omega^{\prime})\Big{\}}^{\frac{1}{\beta}}\mu(\text{\rm{d}}y)\text{\rm{d}}u.$
By (2.9) for $\eta=\alpha(\xi)^{k-1}$ and $p=\frac{k}{k-1}$, we obtain $\|\alpha(\xi)\|_{L^{k}(\mathbb{P})}\leq\alpha(\|\xi\|_{L^{k}(\mathbb{P})})$, which together with (2.18) implies
$\int_{\Omega\times\Omega}\bigg{|}(U^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-U^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))$
$\qquad\qquad\qquad\quad-(\tilde{U}^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-\tilde{U}^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))\bigg{|}^{k}\text{\rm{d}}\mathbb{P}(\omega)\text{\rm{d}}\mathbb{P}(\omega^{\prime})$
$\leq 2^{k}\mathbb{E}\alpha(|X_{s}^{y+\varepsilon u\phi(y),\mu}-\tilde{\theta}^{\lambda,\mu}_{s}(y+\varepsilon u\phi(y))|)^{k}$
$\leq 2^{k}\alpha\left((\mathbb{E}|X_{s}^{y+\varepsilon u\phi(y),\mu}-\tilde{\theta}^{\lambda,\mu}_{s}(y+\varepsilon u\phi(y))|^{k})^{\frac{1}{k}}\right)^{k}$
$\leq c_{k}\alpha(\sqrt{s})^{k}$
for some constant $c_{k}>0$. So, it follows from the fact that $D^{E}b_{t}^{(1)}(y,\nu)(z)$ is continuous in $(y,\nu,z)\in\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{R}^{d}$ due to ($B_{2}$), (3.47), (3.45), (3.25) and the dominated convergence theorem that
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{|}v_{s}^{\varepsilon,1}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r-v_{s}^{\varepsilon,1}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)\Big{|}=0.$
Similarly, by (3.3) and (3.39), we have
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{|}v_{s}^{\varepsilon,2}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r-v_{s}^{\varepsilon,2}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)\Big{|}=0.$
Therefore, (3.46) holds, which implies (3.43) as explained before (3.46).
Moreover, by (3.43), (3.34), (2.18), (3.41), (3.3) and the argument leading to (3.21), we obtain from the dominated convergence theorem that
(3.48) $\lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\int_{\mathbb{R}^{d}}\|\mathbb{E}[U_{t}(\cdot,\cdot,X_{t}^{x,\mu})h_{t}(\varepsilon,x)]\|_{\infty}\mu(\text{\rm{d}}x)=0,$
@@
$\lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\int_{\mathbb{R}^{d}}\bigg{\|}\mathbb{E}\bigg{[}U_{t}(\cdot,\cdot,X_{t}^{x,\mu})$